Solve the following equation by completing the square. (Enter solutions from smallest to largest. If there is one solution with a multiplicity of two, enter it in both answer boxes.)
x^2-4x=32
x = ?
x = ?
Please show me the steps to get the answer. Thanks

Answer by jorel1380(3719)

x2-4x=32
x2-4x-32=0
(x-8)(x+4)=0
x=8 or -4..